In 1985,B is 13. A is 43. In what year will B be 2/5 of A's age?
Answer by checkley77(12844)

(13+X)=2/5(43+X)
13+X=86/5+2X/5
X-2X/5=86/5-13
.6X=(86-13*5)/5
.6X=(86-65)/5
.6X=21/5
X=(21/5)/.6
X=21/5*1/.6
X=21/3
X=7 YEARS AFTER 1985 OR 1992 B WILL BE 2/5 THE AGE OF A.
PROOF:
13+7=2/5(43+7)
20=2/5*50
20=100/5
20=20